# HSM: Hierarchical Scene Motifs for Multi-Scale Indoor Scene Generation

### v2

@@ 1 Introduction
Digital 3D indoor scenes are widely used in domains such as gaming, interior design, virtual training, and simulation for embodied AI and robotics. Consequently, the efficient and controllable generation of realistic indoor scenes has been a long-standing research problem. While recent advances have improved scene synthesis through various approaches, most efforts have focused on arranging large furniture pieces, with less attention given to arrangement of smaller objects, such as computer peripherals and place settings. These objects are often treated as an afterthought, placed randomly or based on predefined rules, limiting realism of the generated scenes and downstream applications.
@@
In this work, we introduce Hierarchical Scene Motifs (HSM), a hierarchical approach to generating indoor scenes with densely populated objects. We employ a unified approach to placing objects based on text-explicit and implicit relationships, from room-level furniture arrangements to the arrangement of small objects on furniture. At each level, HSM first identifies valid support regions (i.e., regions in which objects can be placed) and then generates compositional object arrangements within them. This decomposition transforms the scene generation problem into equivalent subproblems of arranging objects on surfaces, while ensuring that the generated scene aligns with user intent.
@@ 2 Related Work
Indoor scene generation. Early approaches primarily relied on rule-based  [35, 7, 78, 48, 8, 55] or data-driven  [31, 15, 6, 16, 60, 25, 19, 54, 33, 90] methods. The advent of deep learning led to a shift towards learning-based methods [70, 58, 37, 92, 53, 72, 11, 51, 50, 68, 39, 26, 79, 49, 86, 87, 64, 74, 76, 83, 82, 38]. These methods take various forms of input, including text descriptions, scene graphs, and images, to generate 3D scenes accordingly. Of particular interest are methods that generate scenes from text  [68, 39, 83, 82], with recent works increasingly incorporating large language models (LLMs) into the process  [81, 5, 20, 27, 73, 42, 80, 69, 3, 77, 75, 63, 46, 43, 65, 45, 10, 12, 14, 41, 4, 62]. While these methods enable text-conditioned scene generation, most focus solely on arranging large furniture, neglecting the small objects that are ubiquitous in indoor environments. More recent works have begun integrating small objects into the scene generation process  [81, 20, 75, 73, 42, 89, 91, 45, 12, 41, 52, 44, 56, 29, 93, 23]. Additionally, some methods attempt to place small objects in existing scenes  [47, 85, 28, 1]. However, these methods often treat small objects in a simplified or specialized manner, such as using random placement. This limits both the diversity and controllability of their arrangements. Our work introduces a unified hierarchical framework for scene generation that conditions object placement on precise language descriptions, capturing object relationships at all scales.
@@ 3 Method
Given a text description $T$ of an indoor scene and an optional room boundary as a list of vertices as input, HSM generates the scene iteratively through a unified hierarchical framework, as shown in Fig. 2. It first uses a VLM to extract a room type and decompose $T$ into a list of required objects at each scale (Sec. 3.2). The scene is then constructed through three key steps at each hierarchical level: support region extraction (Sec. 3.3), scene motif generation (Sec. 3.4), and layout optimization (Sec. 3.5). These steps are iteratively applied to an initially empty scene, first placing room-level furniture and then arranging small objects on furniture surfaces. After placing the initial objects, if the occupancy is below a threshold $t_{\text{occ}}$ and the number of iterations below $t_{\text{iter}}$, a VLM is prompted for additional objects to add to the scene beyond the input description, conditioned based on the predicted room type. Throughout this process, HSM leverages a library of learned motifs to compose complex spatial relationships as scene motifs (Sec. 3.1), which are instantiated to produce physically valid object arrangements in the scene.
@@ 3.1 Scene Motifs
As in SMC, motifs are implemented as visual programs that encode spatial relationships between objects using programmatic constructs. Each motif is defined as a Python function that takes specific arguments to generate object arrangements when executed. For example, a stack motif may be implemented as a function that iteratively stacks objects vertically using a for loop, with parameters for count and object type (e.g., stack(4, book) to generate “a stack of four books”). HSM assumes the existence of a library of learned motifs, which it dynamically composes to create scene motifs. See the supplement for a complete list of motifs used and details on the learning process.
@@ 3.2 Input Description Requirement Decomposition
Given the input text description, HSM first uses a VLM to extract the room type and an initial list of objects to include in the scene. Objects are grouped by the VLM according to their supporting architectural elements or objects. This breaks the generation task into smaller, more manageable subproblems that can be addressed within a unified framework while minimizing the risk of losing details in subsequent steps (see Fig. 2). For each object, we extract its quantity, appearance, and dimensions from the input description, inferring plausible values when unspecified. If the input lacks a room boundary, we prompt the VLM to generate one including door, window placement and room height. See the supplement for the VLM prompts used.
@@ 3.3 Support Region Extraction
While room-level support regions can be directly extracted from the room boundary, furniture-level support regions require more complex analysis due to the geometric intricacies of furniture surfaces (e.g., multi-level shelves). We first identify horizontal and vertical surfaces on a furniture mesh with $n$ vertices and $p$ triangle faces $F\in\{1,\ldots,n\}^{p\times 3}$, parameterizing each surface as a mesh subset of $F$. Inspired by [32], we cluster $F$ into planar surfaces by seeding each cluster $c_{j}$ with the largest unclustered $f_{j_{0}}$ by area and adding $f_{i}$ to $c_{j}=\{j_{k}\mid k=1,\ldots,\lvert c_{j}\rvert\}$ if its normal is in the same direction as the normals of the first triangle and adjacent triangle. Formally, we check if $n_{i}\cdot n_{j_{0}}\geq t_{\text{norm}}$ and $n_{i}\cdot n_{j_{\text{adj}}}\geq t_{\text{adj}}$, where $n_{i}$ is the normal of $f_{i}$ and $f_{j_{\text{adj}}}$ shares an edge with $f_{i}$. While building up a cluster, we traverse adjacent faces $f_{i}$ by area (from largest to smallest).
We identify horizontal and vertical surfaces by fitting a plane to each $c_{i}$ and thresholding the vertical component of the normal. To ensure functional utility, we compute $h_{i}$ between horizontal surfaces and discard those which have low clearance ($h_{i}<t_{\text{clear}}$). For top surfaces without a ceiling, we assign a default clearance of $h_{\text{top}}$, in order to put a reasonable limit on the size of small objects placed on top of furniture. Finally, horizontal surfaces are split by intersecting vertical surfaces thicker than $t_{\text{seg}}$, ensuring continuity within each region. Fig. 4 illustrates this process for a shelf unit. See the supplement for further implementation details.
@@ 3.4 Scene Motif Generation
We populate $S$ by generating scene motifs $M$. Given the set of objects $O_{i}$ selected for each $m_{i}\in M$ and the overall room description, HSM prompts a VLM to produce a natural language description of object layout. Using this description, we further prompt the VLM to group objects based on their spatial and functional relationships into scene motifs. For example, given the description, “a room with six chairs around a table and a sofa facing a TV,” in the floor support region, the system separately groups the chairs around the table and the sofa facing the TV.
@@
The scene motif is instantiated by creating the primary motif and progressively traversing the hierarchy to model additional objects. We first retrieve a suitable 3D mesh for each object from a database based on its category, appearance description, and dimensions (see the supplement for details on the retrieval process). For each motif, we use a VLM to identify a corresponding program from the library of learned motifs and generate an appropriate function call instantiating the program. This function is then executed to generate an object arrangement using the retrieved meshes. The resulting arrangement is subsequently used as input for the next motif in the hierarchy, iteratively constructing the scene motif until all objects are placed. To ensure physical plausibility between objects, we apply the same spatial optimization used in SMC—resolve collisions, move objects closer, and simulate gravity—while respecting the hierarchy. Motifs are treated as single units during optimization.
@@ 3.5 Layout Optimization
The final step in populating a support region is placing the generated scene motifs within it. This is achieved through a two-stage process. In the first stage, we provide an orthogonal projection of the support region, the bounding boxes of the scene motifs, and the sub-scene description to a VLM, prompting it to suggest initial placements and determine whether the scene motifs should be wall-aligned. The VLM is instructed to consider both explicit constraints from the input description and implicit constraints derived from common arrangements and usage patterns to generate a reasonable initial layout. The second stage refines the layout using an optimization solver. Inspired by Holodeck [81], we employ a grid-based depth-first search (DFS) solver. Starting with the scene motif that has the largest footprint, the solver iteratively refines placements by enforcing the following constraints: 1) scene motifs must be placed within the support region, 2) wall-aligned scene motifs must have their back against a wall and face into the room, and 3) there should be no overlap between scene motifs. The solver returns the first valid layout or otherwise the initial placement if a set time limit is exceeded. Finally, we apply a scene-level spatial optimizer to refine placements by eliminating mesh collisions and ensuring valid support. For each scene motif, we resolve collisions and support issues using relationship-aware rules and raycasting, making minimal adjustments while preserving hierarchy and layout. We provide the details in the supplement.
@@ 4.1 Scene Generation
Input Text Descriptions. For all methods, we use the 100 text descriptions from SceneEval-100, a dataset introduced by SceneEval [67], a recent framework for evaluating text-to-indoor scene generation methods. The descriptions vary in complexity (40 easy, 40 medium, 20 hard) and encompass a wide range of scene types and object arrangements. Each scene description is paired with human-annotated ground truth scene properties (e.g., object counts, attributes, and spatial relationships) to assess how well the text match of the generated scenes.
@@
We also conduct a user perceptual study to compare generated scenes from HSM and Holodeck. We select Holodeck as the baseline for comparison due to its strong generation performance and ability to place small objects. The study consists of two parts: scene-level evaluation and small object-level evaluation. At the scene level, we randomly select 25 scene pairs generated by each method for the same input description. Given top-down renderings, participants assess which scene better aligns with the input text descriptions (Fidelity) and exhibits more physically realistic object placements (Plausibility). At the small object level, we randomly select 30 pairs of furniture populated with small objects and provide close-up renderings. Each pair originates from the same scene description, and participants evaluate the placements based on fidelity and plausibility. The study was conducted with 25 participants, and we report the percentage of participant preferences for each method across both evaluation criteria and levels. See the supplement for the user study instructions.
In addition, we further evaluate alignment between generated scenes and the input descriptions by rendering top-down views and computing text-image similarity using BLIP-2 [36], Long-CLIP [88] and VQAScore [40]. Similarly, we report alignment between small object arrangements and their corresponding text descriptions using the same metrics and the set of 30 populated furniture from the user study. For methods that do not generate small objects, we select the best-matching furniture available for evaluation. The score is set to 0 if no matching furniture of the same category is found.
@@ 4.2 Support Region Extraction
As support region extraction from 3D furniture meshes is a key component of our approach and an essential step for placing small objects, we evaluate this step in isolation. To this end, we manually annotate 100 furniture items (e.g., tables, shelves, sofas) from the HSSD-200 dataset [34] with ground truth support regions. Each annotated region is represented by a set of faces defining the support surface and its height clearance. A single furniture piece may contain multiple support regions of varying shapes and sizes (e.g., as in Fig. 4). The 100 objects we annotated have a total of 529 regions, with an average of five regions per object. See the supplement for details on the annotation process.
@@ 5 Results
Figure 7: Support region extraction examples. Colored boxes are extracted support regions that can be populated with objects.
@@ 5.1 Scene Generation
Fig. 5 presents generated scenes from HSM and the baselines, while Tab. 1 reports the quantitative evaluation results. HSM outperforms the baselines as measured by text-image score and SceneEval fidelity metrics, demonstrating better alignment between generated scenes and input descriptions. Qualitative results further highlight HSM ’s ability to capture precise requirements specified in the input. For example, HSM is the only method that accurately generates a small table next to a loveseat with a coffee table in front in the first row, and two wall sconces flanking a mirror in the third row. In contrast, LayoutGPT frequently places objects outside the room. InstructScene and LayoutVLM also place objects outside room boundaries, and exhibit limited object variety and layout diversity.
Tab. 3 shows that HSM outperform prior work on small object placements. Results from our user study (Tab. 2) further demonstrate that HSM is preferred at both the scene and small object levels in fidelity and plausibility. While Holodeck can generate small objects, it only considers the top surfaces of furniture items, leaving interiors unrealistically empty (as shown in Fig. 6). In contrast, HSM’s support region analysis identifies valid placement regions across all object surfaces, enabling denser and more realistic small object arrangements. By leveraging a hierarchical approach, HSM produces more structured and realistic object placements across all scales within a unified framework. See the supplement for more qualitative results.
@@ 5.2 Support Region Extraction
Fig. 7 shows examples of extracted support regions. HSM achieves an average IoU of $61.20\%$ and F1-score of $52.00\%$ against ground truth annotations. In contrast, the baseline that considers only top surfaces performs significantly worse, with an IoU of $31.12\%$ and an F1-score of $15.82\%$. This demonstrates that HSM more effectively identifies valid support regions within furniture items, which is essential for accurate small object placement in scenes.

### v3

@@ 1 Introduction
Digital 3D indoor scenes are widely used in domains such as gaming, interior design, virtual training, and simulation for embodied AI and robotics. Consequently, efficient and controllable generation of realistic indoor scenes has been a long-standing research problem. While recent advances have improved scene synthesis through various approaches, most efforts have focused on arranging large furniture pieces, with less attention given to arrangement of smaller objects, such as computer peripherals and place settings. These objects are often treated as an afterthought, placed randomly or based on predefined rules, limiting realism of the generated scenes and downstream applications.
@@
In this work, we introduce Hierarchical Scene Motifs (HSM), a hierarchical approach to generating indoor scenes with densely populated objects. We employ a unified approach to placing objects, guided by both text-explicit and implicit relationships. This spans scales from room-level furniture layouts to the placement of small objects on furniture. At each level, HSM first identifies valid support regions (i.e., regions in which objects can be placed) and then generates compositional object arrangements within them. This transforms the scene generation problem into equivalent subproblems of arranging objects on surfaces, while ensuring that the generated scene aligns with user intent.
@@ 2 Related Work
Indoor scene generation. Early approaches primarily relied on rule-based  [36, 8, 80, 49, 9, 56] or data-driven  [32, 16, 7, 17, 61, 26, 20, 55, 34, 92] methods. The advent of deep learning led to a shift towards learning-based methods [71, 59, 38, 94, 54, 74, 12, 52, 51, 69, 40, 27, 81, 50, 88, 89, 65, 76, 78, 85, 84, 39]. These methods take various forms of input, including text descriptions, scene graphs, and images, to generate 3D scenes. Of particular interest are methods that generate scenes from text  [69, 40, 85, 84], with recent works increasingly incorporating large language models (LLMs) into the process  [83, 6, 21, 28, 75, 43, 82, 70, 3, 79, 77, 64, 47, 44, 66, 46, 11, 13, 15, 42, 5, 63]. While these methods enable text-conditioned scene generation, most focus solely on arranging large furniture, neglecting the small objects that are ubiquitous in indoor environments. More recent works have begun integrating small objects into the scene generation process  [83, 21, 77, 75, 43, 91, 93, 46, 13, 42, 53, 45, 57, 30, 95, 24]. Additionally, some methods attempt to place small objects in existing scenes  [48, 87, 29, 1]. However, these methods often treat small objects in a simplified or specialized manner, such as using random placement. This limits both the diversity and controllability of their arrangements. Our work introduces a unified hierarchical framework for scene generation that conditions object placement on precise language descriptions, capturing object relationships at all scales.
@@ 3 Method
Given a text description $T$ of an indoor scene and an optional room boundary as a list of vertices as input, HSM generates the scene iteratively through a unified hierarchical framework, as shown in Fig. 2. It first uses a vision language models (VLM) to extract a room type and decompose $T$ into a list of required objects at each scale (Sec. 3.2). The scene is then constructed through three key steps at each hierarchical level: support region extraction (Sec. 3.3), scene motif generation (Sec. 3.4), and layout optimization (Sec. 3.5). These steps are iteratively applied to an initially empty scene, first placing room-level furniture and then arranging small objects on furniture surfaces. After placing the initial objects, if the occupancy is below a threshold $t_{\text{occ}}$ and the number of iterations below $t_{\text{iter}}$, a VLM is prompted for additional objects to add to the scene beyond the input description, conditioned based on the predicted room type. Throughout this process, HSM leverages a library of learned motifs to compose complex spatial relationships as scene motifs (Sec. 3.1), which are instantiated to produce physically valid object arrangements in the scene.
@@ 3.1 Scene Motifs
As in SMC, motifs are implemented as visual programs that encode spatial relationships between objects using programmatic constructs. Each motif is defined as a Python function that takes specific arguments to generate object arrangements when executed. For example, a stack motif may be implemented as a function that iteratively stacks objects vertically using a for loop, with parameters for count and object type (e.g., stack(4, book) to generate “a stack of four books”). HSM assumes a pre-learned library of motifs exists, and uses it to dynamically composes scene motifs. See Sec. A.1 for more details.
@@ 3.2 Input Description Requirement Decomposition
Given the input text description, HSM first uses a VLM to extract the room type and an initial list of objects to include in the scene. Objects are grouped by the VLM according to their supporting architectural elements or objects. This breaks the generation task into smaller, more manageable subproblems that can be addressed within a unified framework while minimizing the risk of losing details in subsequent steps (see Fig. 2). For each object, we extract its quantity, appearance, and dimensions from the input description, inferring plausible values when unspecified. If the input lacks a room boundary, we prompt the VLM to generate one including door, window placement and room height. See Appendix D for the VLM prompts used.
@@ 3.3 Support Region Extraction
While room-level support regions can be extracted from the room boundary, furniture-level support regions require more complex analysis due to the geometric intricacies of furniture surfaces (e.g., multi-level shelves). We first identify horizontal and vertical surfaces on a furniture mesh with $n$ vertices and $p$ triangle faces $F\in\{1,\ldots,n\}^{p\times 3}$, parameterizing each surface as a mesh subset of $F$. Inspired by [33], we cluster $F$ into planar surfaces by seeding each cluster $c_{j}$ with the largest unclustered $f_{j_{0}}$ by area and adding $f_{i}$ to $c_{j}=\{j_{k}\mid k=1,\ldots,\lvert c_{j}\rvert\}$ if its normal is in the same direction as the normals of the first triangle and adjacent triangle. Formally, we check if $n_{i}\cdot n_{j_{0}}\geq t_{\text{norm}}$ and $n_{i}\cdot n_{j_{\text{adj}}}\geq t_{\text{adj}}$, where $n_{i}$ is the normal of $f_{i}$ and $f_{j_{\text{adj}}}$ shares an edge with $f_{i}$. While building up a cluster, we traverse adjacent faces $f_{i}$ by area (from largest to smallest).
We identify horizontal and vertical surfaces by fitting a plane to each $c_{i}$ and thresholding the vertical component of the normal. To ensure functional utility, we compute $h_{i}$ between horizontal surfaces and discard those which have low clearance ($h_{i}<t_{\text{clear}}$). For top surfaces without a ceiling, we assign a default clearance of $h_{\text{top}}$, in order to put a reasonable limit on the size of small objects placed on top of furniture. Finally, horizontal surfaces are split by intersecting vertical surfaces thicker than $t_{\text{seg}}$, ensuring continuity within each region. Fig. 4 illustrates this process for a shelf unit. See Sec. A.2 for implementation details.
@@ 3.4 Scene Motif Generation
We populate $S$ by generating scene motifs $M$. Given the objects $O_{i}$ selected for each $m_{i}\in M$ and the overall room description, HSM prompts a VLM to produce a natural language description of the object layout. The VLM then uses this description to group objects into $M$ based on spatial and functional relationships. For example, given the description, “a room with six chairs around a table and a sofa facing a TV,” in the floor support region, the system groups the chairs around the table and the sofa facing the TV.
@@
The scene motif is instantiated by creating the primary motif and progressively traversing the hierarchy to model additional objects. We first retrieve a suitable 3D mesh for each object from a database based on its category, appearance description, and dimensions (see Sec. A.4 for details on the retrieval process). For each motif, we use a VLM to select a matching program from a pre-learned library of motifs and generate an appropriate function call to instantiate it. This function is then executed to generate an object arrangement using the retrieved meshes. The resulting arrangement is subsequently used as input for the next motif in the hierarchy, iteratively constructing the scene motif until all objects are placed. To ensure physical plausibility between objects, we apply the same spatial optimization used in SMC—resolve collisions, move objects closer, and simulate gravity—while respecting the hierarchy. Motifs are treated as single units during optimization.
@@ 3.5 Layout Optimization
The final step in populating a support region is placing the generated scene motifs within it. This is achieved through a three-stage process. In the first stage, we provide an orthogonal projection of the support region, the bounding boxes of the scene motifs, and the sub-scene description to a VLM, prompting it to suggest initial placements and determine whether the scene motifs should be wall-aligned. The VLM is instructed to consider both explicit constraints from the input description and implicit constraints derived from common arrangements and usage patterns to generate a reasonable initial layout. The second stage refines the layout using an optimization solver. Inspired by Holodeck [83], we employ a grid-based depth-first search (DFS) solver. Starting with the scene motif that has the largest footprint, the solver iteratively refines placements by enforcing the following constraints: 1) scene motifs must be placed within the support region, 2) wall-aligned scene motifs must have their back against a wall and face into the room, and 3) there should be no overlap between scene motifs. The solver returns the first valid layout or otherwise the initial placement if a set time limit is exceeded. Finally, we apply a scene-level spatial optimizer to refine placements by eliminating mesh collisions and ensuring valid support. For each scene motif, we resolve collisions and support issues using relationship-aware rules and raycasting, making minimal adjustments while preserving hierarchy and layout. We provide the details for the layout optimization in Sec. A.3.
@@ 4.1 Scene Generation
Input Text Descriptions. For all methods, we use the first 100 text descriptions from SceneEval-500, a dataset introduced by SceneEval [68], a recent framework for evaluating text-to-indoor scene generation methods. The descriptions vary in complexity (40 easy, 40 medium, 20 hard) and encompass a wide range of scene types and object arrangements. Each scene description is paired with human-annotated ground truth scene properties (e.g., object counts, attributes, and spatial relationships) to assess how well the text match the generated scenes.
@@
We also conduct a user perceptual study to compare generated scenes from HSM and Holodeck. We select Holodeck as the baseline for comparison due to its strong generation performance and ability to place small objects. The study consists of two parts: scene-level evaluation and small object-level evaluation. At the scene level, we randomly select 25 scene pairs generated by each method for the same input description. Given top-down renderings, participants assess which scene better aligns with the input text descriptions (Fidelity) and exhibits more physically realistic object placements (Plausibility). At the small object level, we randomly select 30 pairs of furniture populated with small objects and provide close-up renderings. Each pair originates from the same scene description, and participants evaluate the placements based on fidelity and plausibility. The study was conducted with 25 participants, and we report the percentage of participant preferences for each method across both evaluation criteria and levels. See Appendix C for the user study instructions.
We also evaluate alignment between generated scenes and input descriptions by rendering top-down views and computing text-image similarity using BLIP-2 [37], Long-CLIP [90] and VQAScore [41]. Similarly, we report alignment between small object arrangements and the corresponding text descriptions using the same metrics and the set of 30 populated furniture from the user study. For methods that do not generate small objects, we select the best-matching furniture available for evaluation. The score is set to 0 if no matching furniture of the same category is found.
@@ 4.2 Support Region Extraction
As support region extraction from 3D furniture meshes is a key component of our approach and an essential step for placing small objects, we evaluate this step in isolation. To this end, we manually annotate 100 furniture items (e.g., tables, shelves, sofas) from the HSSD-200 dataset [35] with ground truth support regions. Each annotated region is represented by a set of faces defining the support surface and its height clearance. A single furniture piece may contain multiple support regions of varying shapes and sizes (e.g., as in Fig. 4). The 100 objects we annotated have a total of 529 regions, with an average of five regions per object. See the Appendix B for details on the annotation process.
@@ 5 Results
Figure 7: Support region extraction examples. Colored boxes are extracted support regions for accurately placing smaller items.
@@ 5.1 Scene Generation
Fig. 5 presents generated scenes from HSM and the baselines, while Tab. 1 reports the quantitative evaluation results. HSM outperforms the baselines as measured by text-image score and SceneEval fidelity metrics, demonstrating better alignment between generated scenes and input descriptions. Qualitative results further highlight HSM ’s ability to capture precise requirements specified in the input. For example, HSM is the only method that accurately generates a small table next to a loveseat with a coffee table in front in the first row, and two wall sconces flanking a mirror in the third row. In contrast, LayoutGPT frequently places objects outside the room. InstructScene and LayoutVLM also place objects outside room boundaries, and exhibit limited object variety and layout diversity. Additional analyses are in the supplement: computational cost and runtime breakdown (Sec. A.6), breakdown of results by difficulty (Tab. 5), and evaluation of HSM using an open-source VLM (Sec. A.7).
Tab. 3 shows that HSM outperform prior work on small object placements. Results from our user study (Tab. 2) further demonstrate that HSM is preferred at both the scene and small object levels in fidelity and plausibility. While Holodeck can generate small objects, it only considers the top surfaces of furniture items, leaving interiors unrealistically empty (as shown in Fig. 6). In contrast, HSM’s support region analysis identifies valid placement regions across all object surfaces, enabling denser and more realistic small object arrangements. By leveraging a hierarchical approach, HSM produces more structured and realistic object placements across all scales within a unified framework. See Fig. 10 and  Fig. 11 for more qualitative results.
@@ 5.2 Support Region Extraction
Fig. 7 shows examples of extracted support regions. HSM achieves an average IoU of $60.27\%$ and F1-score of $48.54\%$ against ground truth annotations. In contrast, the baseline that considers only top surfaces performs significantly worse, with an IoU of $32.54\%$ and an F1-score of $16.80\%$.
This demonstrates that HSM more effectively identifies valid support regions within furniture items, which is essential for accurate small object placement in scenes.